This is a problem for exponential decay. I am just not understanding it.
A radioactive substance has a half-life of 2 years. There is 36 g to begin with. Write and equation for y=mass left after x=half-years.
Then solve for 100 years. (how much of the 36 g will be left?)

Answer by stanbon(57347)

A radioactive substance has a half-life of 2 years.
There is 36 g to begin with.
Write and equation for y=mass left after x=half-years.
Then solve for 100 years. (how much of the 36 g will be left?)
---------------
A(t) = Ao(1/2)^(t/2)
Comment: That t/2 exponent means you should multiply by 1/2 every 2-years.
-----------------
A(100) = 36(1/2)^(100/2)
A(100) = 36(1/2)^50
A(100) = 36*0.00000000000000088817842
A(100) = 0.0000000000000319744...grams
Or
A(100) = 3.197 x 10^-14 grams
==============
============================
Cheers,
Stan H